# Sea-Land-Cloud Segmentation in Satellite Hyperspectral Imagery by Deep Learning

### v2

@@ Appendix B Data Dimensions Arrangement
On the one hand, in the case of 1D-CNN or 1D ML models, we transform each HS image into a 2D array resembling tabular data. Therefore, the models take inputs formatted as BATCH SIZE $\times$ CHANNELS, where BATCH SIZE indicates the number of image pixels for which predictions are to be inferred. 1D-CNN models perform 1D convolutions separately on each pixel across its channels. On the other hand, for 2D-CNNs, optimizing memory and computational resources during training and inference is even more essential, and to help achieving this, we divide each HS image datacube into smaller cubes, often referred as 3D patches or tiles. Padding extends the image borders to retain the integrity of edge information, and the padded borders can be eliminated by cropping the extended spatial dimensions in the segmented image. After padding, we proceed with non-overlapping patching. Unlike overlapping patching, this approach guarantees that each pixel is processed only once to reduce computational complexity. The input format for the 2D-CNNs is BATCH SIZE $\times$ PATCH SIZE $\times$ PATCH SIZE $\times$ CHANNELS, where BATCH SIZE refers in this case to the number of 3D patches. We find that a PATCH SIZE of 48 is suitable for our experiments as it aligns with the encoder-decoder networks, which repeatedly decrease the spatial dimensions by a factor of two via pooling. Consequently, selecting a PATCH SIZE of 48 ensures that the spatial dimensions subjected to pooling remain integer as the network’s depth increases while reducing memory and computational resources in forward propagation.

### v3

@@ Appendix B Data Dimensions Arrangement
Arranging the dimensions of the data is key before model training and inference, as models expect inputs in specific shapes. For 1D-CNN or 1D ML models, we transform each HS 3D data cube into a 2D array resembling tabular data. The models take inputs of the form BATCH SIZE $\times$ CHANNELS, where BATCH SIZE is the number of image pixels for which predictions are to be inferred. 1D-CNNs perform 1D convolutions separately on each pixel across its channels. For 2D-CNNs, optimizing memory and computational resources during training and inference is even more essential, and to help achieve this, we divide each data cube into smaller cubes, often referred as 3D patches. Padding extends the image borders to retain the integrity of edge information, and the padded borders can be eliminated in the final segmented image by cropping the extended spatial dimensions. After data padding, we proceed with non-overlapping patching. Unlike overlapping patching, this approach guarantees that each pixel is processed only once to reduce computational complexity. The input format for the 2D-CNNs is BATCH SIZE $\times$ PATCH SIZE $\times$ PATCH SIZE $\times$ CHANNELS, where BATCH SIZE is the number of 3D patches. We find that a PATCH SIZE of 48 is suitable for our experiments as it aligns with the encoder-decoder networks, which repeatedly decrease the spatial dimensions by a factor of 2 via pooling. Therefore, selecting a small PATCH SIZE of 48 ensures that the spatial dimensions subjected to pooling remain integer as the network’s depth increases. This is also convenient for reducing computational complexity in orbit, as opposed to processing larger patches.